OpenStudy (anonymous):

A chessboard is made of 64 small squares. A single square on a chessboard has an area of 7 cm2. How long is the side of the entire board, rounded to the nearest tenth of a centimeter?

OpenStudy (anonymous):

if the area of a single square is \(7\) then the side of the single square is \(\sqrt{7}\) multiply by the number of squares per side

OpenStudy (anonymous):

I got 4096..

OpenStudy (anonymous):

what did you multiply \(\sqrt{7}\) by ?

OpenStudy (anonymous):

I multiplied eight 8 times

OpenStudy (anonymous):

ok right, the answer is \(8\times \sqrt7\) but that is certainly not that big number you got

OpenStudy (anonymous):

So its 21.2 cm.?

OpenStudy (anonymous):

to the nearest tenth, yes

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw